OpenStudy (anonymous):

a rectangular prism is 5m long, 3 m wide and 6m high. what is the lateral surface area?

OpenStudy (aripotta):

2ab + 2bc + 2ac 2(5x3) + 2(3x6) + 2(5x6) = 30 + 36 + 60 = 126m^2

OpenStudy (anonymous):

the formula of that is 2lw + 2lh + 2wh?

OpenStudy (aripotta):

i believe so, i just googled it lol

OpenStudy (anonymous):

HAHAHAHA